Earley Corporation issued perpetual preferred stock with an 11% annual dividend. The stock currently yields 10%, and its par value is $100. Round your answers to the nearest cent. What is the stock's value? $ Suppose interest rates rise and pull the preferred stock's yield up to 14%. What is its new market value? $
Answer #1

annual dividend=100*11%=11

Stock value=annual dividend/Current yield

a.Stock value=11/0.1=$110

b.Stock value=11/0.14=$78.57(Approx).